OpenStudy (anonymous):

GIllian can spend up to $300 buying both cotton and wool shirts for her club. The cotton shirts cost $15 each and the wool shirts costs $20 each. The number of wool shirts Gillian will buy will be atleast 3, but less than half the number of cotton shirts she buys. Which combination of cotton and wool shirts can he buy? a) 9 cott. & 6 wool b) 10 cott. & 4 wool c) 13 cott. & 4 wool d) 15 cott. & 5 wool

OpenStudy (happy_to_help):

Well lets try and do it with all of your choices a.9 times 15=135 6 times 20 =120 135+120=255 A.255 b.10 times 15=150 4 times 20=80 150+80=230 B.250 c.13 times 15=195 4 times 20=80 195+80=275 C.275 d.15 times 15=225 5 times 20=100 225+100=325 D.325 So the closest to 300 would be C